OpenStudy (anonymous):

If f(x) = x +3/3 , then what is f^–1(x) ?

OpenStudy (anonymous):

3x-3

OpenStudy (anonymous):

\[ We \ write \ like \ this. \ y=(x-3)/3 \ and \ we \ need \ \to\\ get \ x \ rather \ than \ y. \\So \ y=x/3 -1 => x/3=y+1 => x=3y+3So \ f(y)=3y+3. \ Or \ f-1(x)=3x+3\]

OpenStudy (lgbasallote):

if f(x) = \(\large \frac{x+3}{3}\) then you can rewrite it as \[y = \frac{x+3}{3}\] now the first step to finding \(\text f^{-1} (x)\) is interchanging x and y \[x = \frac{y+3}{3}\] now solve for y. start by multiplying both sides by 3 \[3x = y + 3\] now subtract 3 from both sides \[3x - 3 = y\] therefore \[\text f^{-1} (x) = 3x - 3\]